• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Defurne

    Log entry time 05:12:55 on February26,2011

    Entry number 348722

    keyword=Changed the target to LD2 at 5am

    We moved the target several times for Beam Check.
    Then we used LH2 4cm and now LD2 4cm.
    The 15K supply is still at 14.5K.

    FIGURE 1